# A Heptalemma for Quantum Mechanics

### v1

@@ A Heptalemma for Quantum Mechanics
John B. DeBrota & Christian List This version: 1 December 2025. Address for correspondence: Munich Center for Mathematical Philosophy, LMU Munich. Both authors contributed equally to this work. We thank Chris Fuchs, Stephan Hartmann, Levin Hornischer, Beatrice Nettuno, Daniele Oriti, and the participants of a research seminar at the MCMP in November 2025 for helpful conversations and feedback.
@@ 2. Bell’s theorem
Bell’s theorem can be motivated by reference to the famous Einstein-Podolsky-Rosen paradox (EPR) paradox (Einstein et al. 1935). Recall that, in quantum mechanics, it is impossible to measure an individual particle’s position and momentum simultaneously with perfect precision, by Heisenberg’s uncertainty principle. Several of the founders of quantum mechanics, notably Bohr and Heisenberg, took this and other aspects of the new theory to suggest that there is no fact of the matter about a particle’s position or momentum before we perform a measurement. (Perhaps the particle’s position or momentum will become “real” only in the act of measurement or is only a meaningful quantity with respect to the particular experimental arrangement necessary to elicit a value.) Einstein, Podolsky, and Rosen presented a thought experiment that challenges this “no fact” assumption. Suppose we send two maximally entangled particles in opposite directions and let each of them travel very far, in principle even for light-years. The entanglement of the two particles implies that if an observer were to measure the position of one of them, they would be able to predict the outcome of a position measurement of the other with perfect precision. Similarly, if an observer were to measure the momentum of one of the particles, they would be able to predict the outcome of a momentum measurement of the other equally precisely. On the assumption that physics admits no “spooky action at a distance”, especially no transmission of information faster than the speed of light, there is no way any measurement performed on the first particle could instantaneously affect the second particle. It should make no difference to the second particle whether the observer measures the position or the momentum of the first particle. Einstein, Podolsky, and Rosen concluded that the perfect predictability of the measurement outcome for the second particle can only be explained if there are some pre-existing facts about what those measurement outcomes would be. [^2] Here they invoke the principle that “[i]f, without in any way disturbing a system, we can predict with certainty (i.e., with probability equal to unity) the value of a physical quantity, then there exists an element of physical reality corresponding to this physical quantity” (Einstein et al. 1935, p. 777). This goes against Bohr’s and Heisenberg’s “no fact” assumption. Einstein, Podolsky, and Rosen took this to suggest that quantum mechanics is incomplete. It seems that there must be some hidden variables by which the measurement outcomes have been determined in advance. [^3] The EPR paradox is often misrepresented. For an exceptionally clear exposition, we recommend Stacey (2018).
@@
Bell’s theorem is useful not only for diagnosing how quantum mechanics departs from the classical picture of reality, but also for generating a taxonomy of different interpretations of quantum mechanics. These can be classified in terms of which of Bell’s assumptions they give up. Let us briefly illustrate this by reference to a few examples. One class of interpretations gives up the assumption of locality. The de Broglie-Bohm interpretation (Goldstein 2025), for instance, takes this route, by postulating that quantum-mechanical systems behave deterministically based on the global configuration of underlying particle positions. Giving up locality, however, amounts to admitting what Einstein called “spooky action at a distance”. A second class of interpretations gives up measurement independence. So-called superdeterministic interpretations fall into this class. They postulate that there are dependencies between measurement choices and the states of the system that are being measured: different observers’ measurements are correlated due to underlying initial conditions. This kind of interpretation is sometimes characterized as denying that observers have free will with respect to the choice of measurement settings – an implication that some scholars see as a feature and others as a bug. A third class of interpretations, finally, gives up realism as it is defined in Bell’s theorem. This is exemplified by the kind of view associated with Bohr and Heisenberg, often called the Copenhagen interpretation, though there are different variants of it. As already noted, Bohr and Heisenberg thought that there is no fact of the matter about certain properties of a physical system until these are explicitly measured. However, a very diverse range of theories is usually grouped together under the umbrella of the denial of Bell’s realism assumption, and as we will see, there are many distinct ways in which one might relax that assumption. Accordingly, we will show that one can arrive at a more fine-grained taxonomy. [^7] Importantly, of course, the literature already contains several more nuanced no-go results in the vicinity of Bell’s theorem. For example, in a similar spirit to Bell’s theorem, Greenberger et al. (1989) derive a deterministic contradiction between local realism and quantum predictions for three or more particles; Hardy’s paradox (Hardy 1992) shows a contradiction between local realism and quantum predictions using a single set of events; and Leggett’s theorem (Leggett 2003) cuts locality and realism along different seams, ruling out a different class of hidden-variable models if a stronger realism condition is adopted. Subject to a preparation independence assumption, the Pusey-Barrett-Rudolph theorem (Pusey et al. 2012) rules out purely epistemic interpretations of the wavefunction. The Colbeck-Renner theorem (Colbeck and Renner 2011) shows, under assumptions of measurement independence and no-signaling, that no hidden variable extension of quantum theory can improve outcome predictions.
@@ 3. A heptalemma
Interestingly, our no-go result can also be seen as a quantum-mechanical analogue of some recent no-go results from the philosophical literature on first-personal facts. Examples of first-personal facts are “I am having such-and-such experiences” or “I occupy such-and-such position in the world”. First-personal facts are non-objective insofar as they are non-invariant under changes in the subjective perspective from which they hold. Whether it is true that I am having such-and-such experiences obviously depends on who I am and which perspective I occupy. There are at least two no-go results concerning first-personal facts in the recent philosophical literature. One, called a “quadrilemma for theories of consciousness” (List 2025), asserts that realism about first-personal facts is jointly inconsistent with non-solipsism, non-fragmentation, and one world, where these theses are essentially the same as in the present heptalemma. This “quadrilemma” presupposes that first-personal facts are non-relational, and so it implicitly relies on the thesis of non-relationalism as well. [^9] According to realism about first-personal facts, it is not merely the case that “relative to Christian’s perspective, I am having Christian’s experiences”, but “I am having Christian’s experiences” simpliciter. Another related no-go result, due to Fine (2005), asserts a conflict between realism about first-personal facts and three other theses called “first-personal neutrality” (which broadly plays the role of non-solipsism), “absolutism” (broadly equivalent to non-relationalism), and “coherence” (broadly equivalent to non-fragmentation). Fine’s result can be interpreted either as presupposing one world or as treating the combination of non-fragmentation and one world as a single thesis.
@@ 4.1. Rejecting locality
As we have seen, one reaction to Bell’s theorem is to reject locality, i.e., to allow measurements performed in one place to instantaneously bring about effects in a distant place. This, of course, remains a consistent escape route from the heptalemma as well.
@@
A relaxation of locality therefore comes with substantial metaphysical cost. It means admitting what Einstein called “spooky action at a distance”. While he did not live to see Bell’s theorem, it seems likely that Einstein would have entertained almost any other option. Indeed, he argued that rejecting locality imperiled the entire enterprise of science, writing that “[t]he complete suspension of this basic principle would make impossible the idea of (quasi-) closed systems and, thereby, the establishment of empirically testable laws in the sense familiar to us” (Einstein 1948, p. 322, quoted and translated in Howard 1985, p. 188). While many do not share these instincts, locality is certainly not abandoned lightly.
Moreover, any rejection of locality must be done with care as quantum mechanics does not permit signaling, the transmission of information faster than the speed of light. In other words, even if there is instantaneous action at a distance, this physical property cannot be leveraged to enable superluminal communication. The onus seems to be on proponents of non-locality to explain why this should be so.
@@ 4.2. Rejecting measurement independence
A second escape route that we already mentioned in relation to Bell’s theorem is to reject measurement independence. This would permit dependencies between Alice’s and Bob’s measurement choices as well as between the choices of either and the system being measured. However, if Alice’s or Bob’s choices depend on the choices of the other or on the system itself, the sense in which these can be said to be choices at all is called into question. This challenge is not easy to rebut, since the rejection of measurement independence explicitly allows for the possibility that the system in question influences the supposed choices of measurements made upon it. As we also mentioned above, denial of measurement independence is sometimes associated with a denial that experimenters possess free will as far as their choices of measurement settings are concerned. It can of course be contested whether a lack of measurement independence really implies a lack of free will and, if so, whether this speaks against the present escape route. Nonetheless, measurement independence may be rejected in the heptalemma, leaving behind six mutually consistent theses.
Taken to its extreme, one could even reject measurement independence deterministically, such that the choices of Alice and Bob are determined with certainty by the physical state of the system. It is perhaps unsurprising that there are physicists and philosophers who prefer interpretations of this kind, known as superdeterministic interpretations, as this intuition amounts to a supreme dedication to physical determinism. A recent view of this type is the cellular automaton interpretation of ’t Hooft (2016). Here it is imagined that physical reality is fundamentally discrete, local, and deterministic and that this state of affairs is best understood by analogy with models of computation like cellular automata. Everything that happens in reality, including any proposed Bell scenarios and choices of the agents involved, is determined by the initial state of the universe. Another variant of superdeterminism is discussed by Hossenfelder and Palmer (2020) and does not insist on a discrete, computational analogy.
@@ 4.3. Rejecting non-solipsism
While rejecting non-solipsism is a logically possible escape route, it does not appear that anyone seriously entertains it as a reasonable response to the challenges of quantum mechanics. That said, we note that QBism (Fuchs and Schack 2013, Fuchs and Stacey 2018) has been accused of being a kind of solipsistic view. This accusation is derived from QBism’s suggestion that the quantum formalism is a “single-user theory”, that when it is used, it concerns the actions and experiences of a single agent, the user. While this reading of the formalism only requires one observer, QBists are certainly not supposing that there actually is only one. Rather, the quantum formalism, according to QBism, may be adopted by any observer to help her navigate the consequences of her actions. A solipsist could interpret quantum mechanics in a QBist way, but QBism itself does not entail solipsism at all. We say more about QBism in Section 5.
@@ 4.4. Rejecting measurement realism
Both ways of rejecting measurement realism have precedent. Quantum Darwinism (Zurek 2009) is a modern example of a rejection of the first kind. The idea is to try to solve the quantum measurement problem by absorbing the concept of measurement into quantum dynamics. Towards this, it is argued that the phenomenon of decoherence allows one to understand measurement outcomes as quasi-stable aspects of the physical unitary evolution. Zurek writes: “This transition from uncertainty (initial presence of many branches – potential for multiple outcomes) to certainty (once a sufficiently long branch fragment becomes known) accounts for perception of ‘collapse”’ (2009, p. 185). Thus, while we have the perception of stable measurement outcomes, there is no associated fact at the lowest level. We will argue in Section 5.2 that the Everett interpretation (Wallace 2012, Vaidman 2021), also known as “the many-worlds interpretation” or “the relative state formulation”, is best understood, despite these names, as rejecting measurement realism in the second way. Briefly, because all outcomes factually occur, there is no sense in which they occur with well-defined probabilities. Although interpretationally quite far from Everett, one can also make the case that some variants of the Copenhagen interpretation take this route. However, as we elaborate in Section 5.1, Copenhagen is not just one view; while all versions robustly reject Bell’s realism, it does not appear that all versions cleanly reject only one of the four theses that constitute it. Finally, one might also read into QBism an inclination towards rejecting measurement realism in this way, but as we argue in Section 5.3, this is not what the conceptual thrust of QBism most clearly points to.
@@ 4.5. Rejecting non-relationalism
The best-known interpretation which takes this route is aptly named Relational Quantum Mechanics (RQM) (Rovelli 1996, 2025). In RQM, the state of a quantum system is inherently relational, i.e., the state is the relation between observer and system. In this view, there are no stipulations on what counts as an observer – any physical system can have a quantum state relative to any other – but there are no states in a non-relational sense. This commitment to relationalism has been explicitly spelled out by Di Biagio and Rovelli (2021). They write: “Facts happen at every interaction, but they are not absolute: they are relative to the systems involved in the interaction” (p. 29). The authors go on to develop an understanding of what they call stable facts, which can be understood as facts whose relativity can be mostly ignored. Although facts are intrinsically dyadic, of the form $\langle A,B\rangle$, which can be read “$A$ is the case relative to BB”, a stable fact is one that is invariant under changes in the second component; i.e., $\langle A,B\rangle$ is a stable fact if and only if it continues to hold whenever $B$ is replaced by any other $B^{\prime}$. RQM correctly identifies itself as a realist view, moreover, explicitly adhering to the one-non-fragmented-world picture when it describes itself as
@@
Thus RQM retains measurement realism, non-fragmentation, and one world.
Some other interpretations also appear to reject non-relationalism, but arguably none are as thoroughgoing in their rejection of non-relationalism as RQM is. For example, Brukner (2018, p. 7) notes that a possible consequence of his no-go result is that “there cannot be facts of the world per se, but only relative to an observer.” Healey’s pragmatic interpretation (Healey 2012) takes a different route in considering measurement outcomes to be non-relational, yet quantum states are objective and agent-relative.

### v2

@@ A Heptalemma for Quantum Mechanics
John B. DeBrota & Christian List This version: 11 March 2026. Address for correspondence: Munich Center for Mathematical Philosophy, LMU Munich. Both authors contributed equally to this work. We thank Chris Fuchs, Stephan Hartmann, Levin Hornischer, Beatrice Nettuno, Hrvoje Nikolić, Daniele Oriti, Huw Price, Simon Saunders, Blake Stacey, Ken Wharton, and the participants of a research seminar at the MCMP in November 2025 for helpful conversations and feedback.
@@ 2. Bell’s theorem
Bell’s theorem can be motivated by reference to the famous Einstein-Podolsky-Rosen paradox (EPR) paradox (Einstein et al. 1935). Recall that, in quantum mechanics, it is impossible to measure an individual particle’s position and momentum simultaneously with perfect precision, by Heisenberg’s uncertainty principle. Several of the founders of quantum mechanics, notably Bohr and Heisenberg, took this and other aspects of the new theory to suggest that there is no fact of the matter about a particle’s position or momentum before or unless we perform a measurement. (Perhaps the particle’s position or momentum will become “real” only in the act of measurement or is only a meaningful quantity with respect to the particular experimental arrangement necessary to elicit a value.) Einstein, Podolsky, and Rosen presented a thought experiment that challenges this “no fact” assumption. Suppose we send two maximally entangled particles in opposite directions and let each of them travel very far, in principle even for light-years. The entanglement of the two particles implies that if an observer were to measure the position of one of them, they would be able to predict the outcome of a position measurement of the other with perfect precision. Similarly, if an observer were to measure the momentum of one of the particles, they would be able to predict the outcome of a momentum measurement of the other equally precisely. On the assumption that physics admits no “spooky action at a distance”, especially no transmission of information faster than the speed of light, there is no way any measurement performed on the first particle could instantaneously affect the second particle. It should make no difference to the second particle whether the observer measures the position or the momentum of the first particle. Einstein, Podolsky, and Rosen concluded that the perfect predictability of the measurement outcome for the second particle can only be explained if there are some pre-existing facts about what those measurement outcomes would be. [^2] Here they invoke the principle that “[i]f, without in any way disturbing a system, we can predict with certainty (i.e., with probability equal to unity) the value of a physical quantity, then there exists an element of physical reality corresponding to this physical quantity” (Einstein et al. 1935, p. 777). This goes against Bohr’s and Heisenberg’s “no fact” assumption. Einstein, Podolsky, and Rosen took this to suggest that quantum mechanics is incomplete. It seems there must be some hidden variables by which the measurement outcomes have been determined in advance. [^3] The EPR paradox is often misrepresented. For an exceptionally clear exposition, we recommend Stacey (2018).
@@
Bell’s theorem is useful not only for diagnosing how quantum mechanics departs from the classical picture of reality, but also for generating a taxonomy of different interpretations of quantum mechanics. These can be classified in terms of which of Bell’s assumptions they give up. Let us briefly illustrate this by reference to a few examples. One class of interpretations gives up the assumption of locality. The de Broglie-Bohm interpretation (Goldstein 2025), for instance, takes this route, by postulating that quantum-mechanical systems behave deterministically based on the global configuration of underlying particle positions. Giving up locality, however, amounts to admitting what Einstein called “spooky action at a distance”. A second class of interpretations gives up measurement independence. So-called superdeterministic interpretations fall into this class. They postulate that there are dependencies between measurement choices and the states of the system that are being measured: different observers’ measurements are correlated due to underlying initial conditions. This kind of interpretation is sometimes characterized as denying that observers have free will with respect to the choice of measurement settings – an implication that some scholars see as a feature and others as a bug. There is also a different way of relaxing measurement independence, which seeks to preserve the observers’ freedom by postulating retrocausality (see, e.g., Price and Wharton 2015). A third class of interpretations, finally, gives up realism as it is defined in Bell’s theorem. This is exemplified by the kind of view associated with Bohr and Heisenberg, often called the Copenhagen interpretation, though there are different variants of it. As already noted, Bohr and Heisenberg thought that there is no fact of the matter about certain properties of a physical system until these are explicitly measured. However, a very diverse range of theories is usually grouped together under the umbrella of the denial of Bell’s realism assumption, and as we will see, there are many distinct ways in which one might relax that assumption. Accordingly, we will show that one can arrive at a more fine-grained taxonomy. [^7] Importantly, of course, the literature already contains several more nuanced no-go results in the vicinity of Bell’s theorem. For example, in a similar spirit to Bell’s theorem, Greenberger et al. (1989) derive a deterministic contradiction between local realism and quantum predictions for three or more particles; Hardy’s paradox (Hardy 1992) shows a contradiction between local realism and quantum predictions using a single set of events; and Leggett’s theorem (Leggett 2003) cuts locality and realism along different seams, ruling out a different class of hidden-variable models if a stronger realism condition is adopted. Subject to a preparation independence assumption, the Pusey-Barrett-Rudolph theorem (Pusey et al. 2012) rules out purely epistemic interpretations of the wavefunction. The Colbeck-Renner theorem (Colbeck and Renner 2011) shows, under assumptions of measurement independence and no-signaling, that no hidden variable extension of quantum theory can improve outcome predictions.
@@ 3. A heptalemma
Interestingly, our no-go result can also be seen as a quantum-mechanical analogue of some recent no-go results from the philosophical literature on first-personal facts. Examples of first-personal facts are “I am having such-and-such experiences” or “I occupy such-and-such position in the world”. First-personal facts are non-objective insofar as they are non-invariant under changes in the subjective perspective from which they hold. Whether it is true that I am having such-and-such experiences obviously depends on who I am and which perspective I occupy. There are at least two no-go results concerning first-personal facts in the recent philosophical literature. One, called a “quadrilemma for theories of consciousness” (List 2025), asserts that realism about first-personal facts is jointly inconsistent with non-solipsism, non-fragmentation, and one world, where these theses are essentially the same as in the present heptalemma. This “quadrilemma” presupposes that first-personal facts are non-relational, and so it implicitly relies on the thesis of non-relationalism as well. [^9] According to realism about first-personal facts as discussed in List (2025), it is not merely the case that “relative to Christian’s perspective, I am having Christian’s experiences”, but “I am having Christian’s experiences” simpliciter. Another related no-go result, due to Fine (2005), asserts a conflict between realism about first-personal facts and three other theses called “first-personal neutrality” (which broadly plays the role of non-solipsism), “absolutism” (broadly equivalent to non-relationalism), and “coherence” (broadly equivalent to non-fragmentation). Fine’s result can be interpreted either as presupposing one world or as treating the combination of non-fragmentation and one world as a single thesis.
@@ 4.1. Rejecting locality
As we have seen, one reaction to Bell’s theorem is to reject locality, i.e., to allow measurements performed in one place to instantaneously bring about effects in a distant place or to stand in correlations that cannot be accounted for by purely local mechanisms. This, of course, remains a consistent escape route from the heptalemma as well.
@@
A relaxation of locality therefore comes with substantial metaphysical cost. In its most clearcut form, it means admitting what Einstein called “spooky action at a distance”. While he did not live to see Bell’s theorem, it seems likely that Einstein would have entertained almost any other option. Indeed, he argued that rejecting locality imperiled the entire enterprise of science, writing that “[t]he complete suspension of this basic principle would make impossible the idea of (quasi-) closed systems and, thereby, the establishment of empirically testable laws in the sense familiar to us” (Einstein 1948, p. 322, quoted and translated in Howard 1985, p. 188). While many do not share these instincts, locality is certainly not abandoned lightly.
Moreover, any rejection of locality must be done with care as quantum mechanics does not permit signaling, the transmission of information faster than the speed of light. In other words, even if there is non-locality in a physical system, this property cannot be leveraged to enable superluminal communication. We would thus require a form of non-locality that carefully constrains the permissible action at a distance. The onus seems to be on proponents of non-locality to explain why there is non-locality and yet no signalling.
@@ 4.2. Rejecting measurement independence
A second escape route that we already mentioned in relation to Bell’s theorem is to reject measurement independence. This would permit dependencies between Alice’s and Bob’s measurement choices as well as between the choices of either and the system being measured. However, if Alice’s or Bob’s choices depend on the choices of the other or on the system itself, the sense in which these can be said to be choices at all is called into question. This challenge is not easy to rebut, since the rejection of measurement independence seems to allow for the possibility that the system in question influences the supposed choices of measurements made upon it. As we also mentioned above, denial of measurement independence is sometimes associated with a denial that experimenters possess free will as far as their choices of measurement settings are concerned. It can of course be contested whether a lack of measurement independence really implies a lack of free will and, if so, whether this speaks against the present escape route. Nonetheless, measurement independence may be rejected in the heptalemma, leaving behind six mutually consistent theses.
Taken to its extreme, one could even reject measurement independence deterministically, such that the choices of Alice and Bob are determined with certainty by the physical state of the system. It is perhaps unsurprising that there are physicists and philosophers who prefer interpretations of this kind, known as superdeterministic interpretations, as this intuition amounts to a supreme dedication to physical determinism. A recent view of this type is the cellular automaton interpretation of ’t Hooft (2016). Here it is imagined that physical reality is fundamentally discrete, local, and deterministic and that this state of affairs is best understood by analogy with models of computation like cellular automata. Everything that happens in reality, including any proposed Bell scenarios and choices of the agents involved, is determined by the initial state of the universe. Another variant of superdeterminism is discussed by Hossenfelder and Palmer (2020) and does not insist on a discrete, computational analogy. For an approach that rejects measurement independence without postulating determinism, see, for example, the statistical contextual interpretation (Kupczynski 2025).
Superdeterministic approaches give up the idea that measurement settings are exogenous variables whose values can be “freely” chosen by the relevant observers. By contrast, there is also an approach that relaxes measurement independence while still treating measurement settings as exogenous. It is the retrocausal approach defended, for instance, by Price and Wharton (2015); for a survey, see Friederich and Evans (2023). This approach accounts for probabilistic dependencies between the state of the common source and the observers’ measurement settings by allowing the latter to influence the former, rather than the other way around. In effect, this reverses the direction of some of the arrows in the kind of causal graph illustrated in Figure 1 above. This blocks the derivation of the Bell inequalities without abandoning the intuitive idea that the observers’ measurement choices are in some sense free. The cost is the admission of retrocausality.
@@ 4.3. Rejecting non-solipsism
While rejecting non-solipsism is a logically possible escape route, it does not appear that anyone seriously entertains it as a reasonable response to the challenges of quantum mechanics. [^10] Convivial Solipsism (Zwirn 2000) and the solipsistic hidden variable model of Nikolić (2012) both postulate multiple observers and are therefore not true solipsisms under our definition. The latter can be converted into a true solipsism by simply supposing that only one observer exists. As presented, however, objective reality is associated to each conscious observer, and so Nikolić’s model is actually a rejection of the one-world postulate. That said, we note that QBism (Fuchs and Schack 2013, Fuchs and Stacey 2018) has been accused of being a solipsistic view (e.g., Norsen 2014). This accusation is derived from QBism’s suggestion that the quantum formalism is a “single-user theory”, i.e., that when it is used, it concerns the actions and experiences of a single agent, the user. While this reading of the formalism only requires one observer, QBists are certainly not supposing that there actually is only one. Rather, the quantum formalism, according to QBism, may be adopted by any observer to help her navigate the consequences of her actions. A solipsist could interpret quantum mechanics in a QBist way, but QBism itself does not entail solipsism at all. We say more about QBism in Section 5.
@@ 4.4. Rejecting measurement realism
Both ways of rejecting measurement realism have precedent. Quantum Darwinism (Zurek 2009) is a modern example of a rejection of the first kind. The idea is to try to solve the quantum measurement problem by absorbing the concept of measurement into quantum dynamics. Towards this, it is argued that the phenomenon of decoherence allows one to understand measurement outcomes as quasi-stable aspects of the physical unitary evolution. Zurek writes: “This transition from uncertainty (initial presence of many branches – potential for multiple outcomes) to certainty (once a sufficiently long branch fragment becomes known) accounts for perception of ‘collapse”’ (2009, p. 185). Thus, while we have the perception of stable measurement outcomes, there is no associated fact at the lowest level. We will argue in Section 5.2 that the Everett interpretation (Wallace 2012, Vaidman 2021), also known as “the many-worlds interpretation” or “the relative state formulation”, is best understood, despite these names, as rejecting measurement realism in the second way. Briefly, because all outcomes factually occur, there is no sense in which they occur with well-defined probabilities. (In Section 5.2, we will also consider some alternative ways of interpreting Everett.)
Although interpretationally quite far from Everett, one can make the case that some variants of the Copenhagen interpretation reject measurement realism too. However, as we elaborate in Section 5.1, Copenhagen is not just one view; while all versions robustly reject Bell’s realism, it does not appear that all versions cleanly reject only one of the four theses that constitute it. Finally, one might read into QBism an inclination towards rejecting measurement realism, but as we argue in Section 5.3, this is not what the conceptual thrust of QBism most clearly points to.
@@ 4.5. Rejecting non-relationalism
The best-known interpretation which takes this route is aptly named Relational Quantum Mechanics (RQM) (Rovelli 1996, 2025). In RQM, the state of a quantum system is inherently relational, i.e., the state is the relation between observer and system, or even more generally between system and system. In this view, any physical system can have a quantum state relative to any other – so, there are no stipulations on what could count as an “observer” – and there are no states in a non-relational sense. This commitment to relationalism has been explicitly spelled out by Di Biagio and Rovelli (2021). They write: “Facts happen at every interaction, but they are not absolute: they are relative to the systems involved in the interaction” (p. 29). The authors go on to develop an understanding of what they call stable facts, which can be understood as facts whose relativity can be mostly ignored. Although facts are intrinsically dyadic, of the form $\langle A,B\rangle$, which can be read “$A$ is the case relative to BB”, a stable fact is one that is invariant under changes in the second component; i.e., $\langle A,B\rangle$ is a stable fact if and only if it continues to hold whenever $B$ is replaced by any other $B^{\prime}$. RQM correctly identifies itself as a realist view, moreover, explicitly adhering to the one-non-fragmented-world picture when it describes itself as
@@
Thus RQM retains measurement realism, non-fragmentation, and one world. There is also room to define variants of RQM with different degrees of commitment to relationalism (see, e.g., Calosi and Riedel 2024, Stacey 2021).
Some other interpretations appear to reject non-relationalism too, though they differ in what exactly they take outcomes or facts to be relative to, for instance whether the relativization can be to any system or merely to an observer. For example, Brukner (2018, p. 7) notes that a possible consequence of his no-go result is that “there cannot be facts of the world per se, but only relative to an observer.” Healey’s pragmatic interpretation (2012) takes a different route in considering measurement outcomes to be non-relational, yet quantum states are agent-relative.